sinπœƒ+cosπœƒ= 1βˆ’2cos2πœƒ/sin πœƒβˆ’cos πœƒ


1
Expert's answer
2021-03-09T16:27:23-0500

"sin\theta+cos\theta=\frac{1-2cos^2\theta}{sin\theta-cos\theta}"


Rearranging the right hand side(RHS) becomes,


"\frac{-1(2cos^2\theta-1)}{-1(cos \theta-sin\theta)}=\frac{2cos^2\theta-1}{cos\theta-sin\theta}"

From the following trigonometric identity,

"cos^2\theta+sin^2\theta=1"

then, the numerator,

"2cos^2\theta-1=2cos^2\theta-(cos^2\theta+sin^2\theta)"

so,

"2cos^2\theta-1=cos^2\theta-sin^2\theta"

Substituting for the numerator in the RHS,

"sin\theta+cos\theta=\frac{cos^2\theta-sin^2\theta}{cos\theta-sin\theta}"

The numerator of the right hand side is a difference of two squares, thus,

"cos^2\theta-sin^2\theta=(cos\theta+sin\theta)(cos\theta-sin\theta)"

Substituting for the numerator of the RHS,

"sin\theta+cos\theta=\frac{(cos\theta+sin\theta)(cos\theta-sin\theta)}{cos\theta-sin\theta}"


Simplifying the RHS proves that the left hand side is equal to the right hand side, that is,

"sin\theta+cos\theta=cos\theta+sin\theta"


This proves that,

"sin\theta+cos\theta=\frac{1-2cos^2\theta}{sin\theta-cos\theta}"